The gravitational force between two people is 1.50 10^-6N. If person A has a mass of 82.0kg and is 0.600m away from person B person Bs mass is ___kg?
1
Expert's answer
2020-06-08T10:14:00-0400

According to the Newton's gravity law, the force between two bodies with masses "m_1" and "m_2" separated by the distance "r" is the following:


"F = G\dfrac{m_1m_2}{r^2}"

where "G = 6.67\cdot 10^{-11} m^3kg^{-1}s^{-2}" is the gravitational constant. Expressing "m_1" from this formula get:


"m_1 = \dfrac{r^2F}{m_2G}"

Substituting numerical values:


"m_1 = \dfrac{0.6^2\cdot 1.5\cdot 10^{-6}}{82\cdot 6.67\cdot 10^{-11}} \approx 98.7kg"

Answer. m2 = 98.7 kg.